OpenStudy (anonymous):

type 4/5+1/25 into a simplified fraction

OpenStudy (jamesj):

For real? Well 4/5 + 1/25 = 20/25 + 1/25 = 21/25.

OpenStudy (anonymous):

excellent thanks

OpenStudy (anonymous):

im new to this i hate math.... do I give you kudos or anything?

OpenStudy (jamesj):

Click on 'good answer' for me

OpenStudy (jamesj):

it's the etiquette around here :-)